# Bootstrapping and autonomous tuning system for Si/SiGe multi-quantum dot devices

### v3

@@ I Introduction
Quantum dot (QD) qubit devices are commonly formed by electrostatically manipulating the potential energy landscape of a two-dimensional (2D) electron or hole gas (EG or HG). This procedure produces interleaved regions of accumulated and depleted charge density that serve as QD qubits, charge sensors, and charge carrier reservoirs [1, 2, 3, 4, 5]. In recent years, the number of QDs in such devices has risen rapidly [6, 7, 8]. In the near term, it will no longer be possible to tune up QD qubit arrays manually. The field of autotuning has arisen to address this challenge [9, 10].
Modern semiconductor QD qubit devices have two critical features that enhance their performance while simultaneously making autotuning challenging and essential. First, the active regions of the device—the qubit QDs, charge sensors, and reservoirs—contain no doping. All charge carriers—electrons or holes—are induced by gate voltages and no current can flow until gate voltages are applied and carriers are accumulated. At the same time, the charges trapped within the gate-oxide heterostructure of the QD devices make it impossible to know a priori the precise turn-on voltages, i.e., the gate voltage threshold at which carrier accumulation begins. The device drift and hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) introduce additional variability in voltage configuration not only between cooldowns but even between successive initialization steps. Second, QD device chips have multiple current channels that often serve different purposes. Some host an array of QD qubits, while others provide a home for charge sensors. These channels are not separated until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined  [^1] Mathematically, the pinch-off and the turn-on voltage are indistinguishable for a given current curve. However, when measuring a physical gate, a hysteresis can be observed.. The size of the QD introduces additional challenges: the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers. While spatially separated, the channels must remain coupled to ensure proper operation of proximal charge sensors on nearby QDs.
@@
Previous work on autotuning has shown that correctly executed bootstrapping of QD devices sets the stage for the more specialized tuning phases [9]. For example, setting the device topology has been demonstrated for undoped accumulation-mode QD devices [12, 13]. Automated calibration of the charge state has been demonstrated for a modulation-doped QD device [14], an undoped SiMOS QD device [15], and undoped accumulation-mode QD devices [16].
@@
Automated tuning of QD devices from a coldstart has been demonstrated for modulation-doped QD devices with predefined current channels [19, 20, 21] as well as for SiFinFETs and nanowires, which by design have a single current channel [21]. However, all these demonstrations assumed a prior knowledge of the device and a non-negligible device-specific preconfiguration. The first fully autonomous bootstrapping algorithm, intended for initiation on a pristine modulation-doped QD device, was introduced in Ref. [22]. However, the proposed algorithm is platform-specific and assumes the QD device to be functional. Autonomous characterization and bootstrapping for accumulation-mode QD devices have never been attempted.
BATIS addresses a key challenge in bootstrapping the QD devices: the diversity of device architectures and control layouts. All initialization tools proposed to date rely on rigid measurement sequences tailored to specific device designs, limiting their applicability across platforms. BATIS, in contrast, is designed to be platform-agnostic and thus can operate across a wide range of QD devices by dynamically adapting its procedures. While the procedural flow of BATIS is well-defined at a high level, the executable flow—the actual sequence of measurements and operations—is automatically adjusted in real time based on the device’s configuration and control wiring, as specified in a flexible configuration file. In other words, throughout execution, BATIS adjusts which steps to take and what outcomes to expect based on the device’s observed behavior. As tuning progresses, BATIS further tailors subsequent measurements to match the discovered characteristics of the system, ensuring efficient and device-specific calibration without manual intervention.
Importantly, the sample used to demonstrate BATIS is kept at around 1.3 $\mathrm{K}$, which is about 10 times warmer than other typical devices [23, 24]. High temperature operation is critical for practical quantum computer implementations [25]. However, higher temperatures broaden key measurement features, such as Coulomb blockade peaks, necessitating careful fitting procedures to extract the desired parameters [26]. The industry has demonstrated promising efforts in characterizing the performance of industry-manufactured spin qubit devices at 1.6 $\mathrm{K}$ using the cryogenic 300-$\mathrm{mm}$ wafer prober [27, 28]. Autotuning of a single-electron transistor (SET) charge sensor in Si at 1.5 $\mathrm{K}$ has also been demonstrated [29].
@@
This paper is organized as follows: In Sec. II, we motivate the necessity of autotuning from the perspective of trapped charges. The bootstrap tuning procedure is described in detail in Sec. III. The experimental demonstration of BATIS to fully autonomously tune-up an accumulation mode Si/SixGe1-x QD device at 1.3 $\mathrm{K}$ is described in Sec. IV. We conclude with a discussion of future directions in Sec. V.
@@ II Trapped Charge: A Barrier to Robust Control
A primary challenge in automating current channel formation in QD devices stems from trapped charges within the QD array, which introduce substantial variability in the required gate voltages. Charge trapping refers to the immobilization of carriers at unknown locations due to, e.g., structural imperfections and defects on surfaces and interfaces of the heterostructure hosting the QDs. Stray charges can accumulate in any region of the heterostructure, including at buried interfaces, in the oxide layer at the top of the stack, and in particular, at the semiconductor-oxide interface, where there is an extremely high density of interface states [30].
@@
While trapped charges are largely locked into place as the device is cooled down from room temperature, they can nevertheless slowly drift over time, even at $\mathrm{mK}$ temperatures [32]. Moreover, these processes are not well controlled, leading to nonuniform charge accumulation that can vary from one cooldown to the next. As a result, significant and nonuniform voltage shifts may be required to offset the effects of trapped charges, and to achieve uniform gate operation across a device, even for nominally identical gate electrodes within a single device.
@@
The effect of trapped charges can be illustrated by simulating channel formation in a typical QD device. Figures 1(a-b) show a gate-defined, accumulation-mode, undoped QD device with multiple current channels, six designated regions for forming QDs (under the six plunger gates, shown in red), and reservoir gates (shown in blue) to connect these regions to ohmic contacts for current flow. This device exemplifies a commonly used fabrication strategy known as overlapping gates [35, 36].
@@
The screening gates (shown in yellow) form the bottommost layer in the three-layer gate architecture, positioned closest to the gate oxide. Their primary role is to screen electric fields, thereby preventing the unintentional formation of 2DEG or 2DHG regions and electrically isolating the intended transport channels. The second layer includes the reservoir and plunger gates. The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device, where QDs will form in the electrostatic confinement potential. Each plunger gate primarily controls the chemical potential of the QD below it. The final layer consists of barrier gates (shown in gray), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and reservoirs. By carefully adjusting the voltages applied to each of the gates, the QD qubit chip with charge sensors is made functional [2, 3, 4, 5].
@@
Figure 1: (a) False-color scanning electron microscopy (SEM) image of a device nominally identical to one used in this work. The color indicates the gate’s function as a screening (Si, $i=1\dots,4$; yellow), reservoir (Ri, $i=1\dots,5$; blue), barrier (Bi, $i=1\dots,9$; gray), or plunger (Pi, $i=1\dots,6$; red) gate. (b) Higher magnification image with the desired current paths shown by black arrows and the individual barrier and plunger gates labeled. (c-f) Self-consistent Schrödinger-Poisson simulations (density plots), showing typical effects of trapped charge on 1D channel formation in a 2DEG. The simulation results are plotted over the top of the SEM image. (c) Trapped charge can lead to current paths crossing the central screening gate; an autotuner must detect and correct for this problem, as shown in (d). (e) Trapped charge can also lead to breaks in the current paths, another possible error that an autotuner must detect and correct for, as shown in (f). (c-f) Insets show the gate voltages necessary to cause the charge densities. For all simulations, interface trapped charge is chosen, as described in the main text and in the Supplemental Material [37].
@@
To model the effects of trapped charge, we introduce randomly varying charge densities on a tiled grid array at the oxide-semiconductor interface, with the density of each tile selected at random from a normal distribution with the standard deviation corresponding to a shift magnitude of $\Delta V=83$ $\mathrm{mV}$, consistent with experimental observations for silicon QDs [28]. Each tile is $40\times 50$ $\mathrm{nm}$, ensuring that all gates encompass multiple regions of randomly selected charge densities.
@@
Figure 2: BATIS algorithm. The flow chart representation of the procedural flow of BATIS. Beyond manipulating gate voltages, BATIS can also change the device response to gate voltages in-situ by illuminating. Active steps in autotuning are shown by gray rounded rectangles, conditional logic by yellow diamonds, and failure determinations by red stop signs. The green and red arrows indicate the passing or failing of a given step, respectively.
@@
In the case where no random trapped charge is included, the ideal charge densities in the quantum well form coherent channels halfway between the screening gates, as indicated with the arrows in Fig. 1(b). In contrast, Figs. 1(c) and 1(e) show results obtained when randomly distributed trapped charges are included, resulting in current channels that flow beneath screening gates and pinched channels, respectively. Both problems can be resolved by careful adjustment of the various gate voltages to recover the coherent channels shown in Figs. 1(d) and 1(f), respectively.
@@
While misdirected and pinched channels are easy to identify in simulations, they are much more difficult to diagnose in experiments. The issue of current flowing beneath the screening gate and unintentionally connecting multiple channels, as illustrated in Fig. 1(c), is intrinsic to 2D gate-defined systems and does not exist in 1D nanowire and nanotube devices [38, 39]. As such, it is of particular interest in this work. The recovery of misdirected current to form the desired coherent channels requires the simultaneous tuning of multiple gates. The protocols developed in this work are specially designed to address and characterize such 2D challenges.
@@ III BATIS: Design and implementation
BATIS is designed to automate the testing, characterization, and initialization of all gates and the forming of all necessary 1D current channels for linear QD arrays. The procedural flow of the single-button-press algorithm, depicted in Fig. 2, is rooted in the underlying device physics, making it both interpretable and robust, while also enabling adaptability across devices and scalability to arbitrary-length 1D arrays.
@@
The configuration files establish a link between abstract descriptions of device components and the operational language used within BATIS. For instance, the functionality of each gate is automatically inferred from the device architecture provided by the user, enabling BATIS to construct a gate lookup table. When combined with the wire map, this allows BATIS to automatically identify and perform the appropriate specialized measurements by linking the logical gate definitions to the physical measurement instruments. This abstraction layer ensures that BATIS remains agnostic to both the specific device layout and the instrumentation used.
@@
The algorithm follows conditional logic, visually represented with boxes and arrows of different colors in Fig. 2. Each of the six stages in BATIS, indicated in Fig. 2 with gray rounded rectangles, involves at least one optimization of a device measurement or manipulation followed by data analysis to extract characteristics of a selected subset of gates. The measurements typical of each stage are detailed in Fig. 3, with the top row in each panel explaining the measurement, and the middle and bottom rows showing for each step a cartoon representation of the data (in each case, showing success in the left and failure on the right) and an example of measured experimental data, respectively.
@@
The procedural steps in BATIS are indicated with gray arrows. Passing the functionality test for the current stage, marked with yellow diamonds in Fig. 2, initiates the next stage, as indicated with green arrows. Failing a test, indicated with red arrows, results in repeated measurements, device optical illumination to adjust the trapped charges, or bootstrapping termination, depending on the issues revealed by BATIS and the conclusions it draws based on the device physics.
@@
Once completed, BATIS produces a device diagnostic file that stores all information about each gate learned during tuning. When executed successfully, it verifies that all gates are functioning as intended and that isolated regions can be formed in well-defined 1D channels beneath the designated plunger gates. It also preconfigures the QD device for further use. In the following, we explain each state of the algorithm.
Figure 3: Stages of the tuning procedure. The top row in each panel explains the measurements. The center row shows idealized cartoon representations of the key data for each stage of the tuning process, with good and bad data shown on the left and the right, respectively. The bottom row provides examples of experimental data. The yellow arrows in the cartoons indicate the ideal current paths. (a) In the leakage test, no gate should have a low resistance to any other gate or cryostat ground (ohmic contacts are included in the test for the non-accumulated leakage test). The good cartoon matrix shows high-resistance paths on all connections. The bad cartoon matrix shows two gates shorted to each other. The realistic matrix shows three gates shorted to ground. Devices with shorted gates are typically unusable. (b) When accumulating the device globally, the current is expected to rise above zero with the increase of gate voltages, as seen in the good cartoon plot. The bad cartoon plot shows a failure to turn on. A successful turn-on is observed in the experimental data. (c) For pinch-off, it is desirable for the current to drop to the noise floor as the gate voltage is lowered. The bad cartoon demonstrates a failure of the screening or reservoir gate to pinch off. The experimental data show some small oscillations due to, e.g., accidental Coulomb blockade during pinch-off. (d) For channel accumulation, rather than relying on measuring pairs of screening gates, we sweep a single screening gate (S) against all the finger gates corresponding to that current channel, ($\mathcal{F}$). The bad cartoon measurement indicates that at least one finger gate is too far below the threshold to accumulate. (e) Each plunger and barrier gate must pinch off individually. The bad cartoon demonstrates a failure of the finger gate to pinch off. In the experimental data, Coulomb blockade oscillations are often present and prominently observed during pinch-off at this tuning step.

### v4

# Bootstrapping, autonomous testing, and initialization system for Si/SixGe1-x multi-quantum-dot devices

@@ I Introduction
Quantum dot (QD) qubit devices are commonly formed by electrostatically manipulating the potential energy landscape of a two-dimensional (2D) electron gas (2DEG) or hole gas (2DHG). This procedure produces interleaved regions of accumulated and depleted charge density that serve as QD qubits, charge sensors, and charge carrier reservoirs [1, 2, 3, 4, 5]. In recent years, the number of QDs in such devices has risen rapidly [6, 7, 8]. In the near term, it will no longer be possible to tune up QD qubit arrays manually. The field of autotuning has arisen to address this challenge [9, 10].
Modern semiconductor QD qubit devices have two critical features that enhance their performance while simultaneously making autotuning challenging and essential. First, the active regions of the device—the qubit QDs, charge sensors, and reservoirs—contain no doping. All charge carriers—electrons or holes—are induced by gate voltages, and no current can flow until gate voltages are applied and carriers are accumulated. At the same time, the charges trapped within the gate-oxide heterostructure of the QD devices make it impossible to know a priori the precise turn-on voltages, i.e., the gate voltage threshold at which carrier accumulation begins. The device drift and hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) introduce additional variability in voltage configuration not only between cooldowns but even between successive initialization steps. Second, QD device chips have multiple current channels that often serve different purposes. Some host an array of QD qubits, while others provide a home for charge sensors. These channels are not separated until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined  [^1] Mathematically, the pinch-off and the turn-on voltage are indistinguishable for a given current curve. However, when measuring a physical gate, a hysteresis can be observed.. The size of the QD introduces additional challenges: the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers. While spatially separated, the channels must remain coupled to ensure proper operation of proximal charge sensors on nearby QDs.
@@
Previous work on autotuning has shown that correctly executed bootstrapping of QD devices sets the stage for the more specialized tuning phases [9]. For example, setting the device topology has been demonstrated for undoped accumulation-mode QD devices [12, 13]. Automated calibration of the charge state has been demonstrated for a modulation-doped QD device [14], an undoped silicon-metal-oxide-semiconductor QD device [15], and undoped accumulation-mode QD devices [16].
@@
Automated tuning of QD devices from a bootstrapping has been demonstrated for modulation-doped QD devices with predefined current channels [19, 20, 21] as well as for silicon in field-effect transistors and nanowires, which by design have a single current channel [21]. However, all these demonstrations assumed prior knowledge of the device and a non-negligible device-specific preconfiguration. The first fully autonomous bootstrapping algorithm, intended for initiation on a pristine modulation-doped QD device, was introduced in Ref. [22]; however, the proposed algorithm is platform specific and assumes the QD device to be functional. Autonomous characterization and bootstrapping for accumulation-mode QD devices have never been attempted to our knowledge.
BATIS addresses a key challenge in bootstrapping the QD devices: the diversity of device architectures and control layouts. All initialization tools proposed to date rely on rigid measurement sequences tailored to specific device designs, limiting their applicability across platforms. BATIS, in contrast, is designed to be platform agnostic and thus can operate across a wide range of QD devices by dynamically adapting its procedures. While the procedural flow of BATIS is well-defined at a high level, the executable flow—the actual sequence of measurements and operations—is automatically adjusted in real time on the basis of the device’s configuration and control wiring, as specified in a flexible configuration file. In other words, throughout execution, BATIS adjusts which steps to take and what outcomes to expect on the basis of the device’s observed behavior. As tuning progresses, BATIS further tailors subsequent measurements to match the discovered characteristics of the system, ensuring efficient and device-specific calibration without manual intervention.
Importantly, the sample used to demonstrate BATIS is kept at around 1.3 $\mathrm{K}$, which is about 10 times warmer than other typical devices [23, 24]. High-temperature operation is critical for practical quantum computer implementations [25]. However, higher temperatures broaden key measurement features, such as Coulomb blockade peaks, necessitating careful fitting procedures to extract the desired parameters [26]. The industry has demonstrated promising efforts in characterizing the performance of industry-manufactured spin qubit devices at 1.6 $\mathrm{K}$ using the cryogenic 300-$\mathrm{mm}$ wafer prober [27, 28]. Autotuning of a single-electron transistor charge sensor in Si at 1.5 $\mathrm{K}$ has also been demonstrated [29].
@@
This paper is organized as follows: In Sec. II, we provide the motivation for the necessity of autotuning from the perspective of trapped charges. The bootstrap tuning procedure is described in detail in Sec. III. The experimental demonstration of BATIS to fully autonomously tune up an accumulation-mode Si/SixGe1-x QD device at 1.3 $\mathrm{K}$ is described in Sec. IV. We conclude with a discussion of future directions in Sec. V.
@@ II Trapped Charge: A Barrier to Robust Control
A primary challenge in automating current channel formation in QD devices stems from the presence of trapped charges within the QD array, which introduce substantial variability in the required gate voltages. “Charge trapping” refers to the immobilization of carriers at unknown locations due to, for example, structural imperfections and defects on the surfaces and at interfaces of the heterostructure hosting the QDs. Stray charges can accumulate in any region of the heterostructure, including at buried interfaces, in the oxide layer at the top of the stack, and in particular, at the semiconductor-oxide interface, where there is an extremely high density of interface states [30].
@@
While trapped charges are largely locked into place as the device is cooled down from room temperature, they can nevertheless slowly drift over time, even at millikelvin temperatures [32]. Moreover, these processes are not well controlled, leading to nonuniform charge accumulation that can vary from one cooldown to the next. As a result, significant and nonuniform voltage shifts may be required to offset the effects of trapped charges and achieve uniform gate operation across a device, even for nominally identical gate electrodes within a single device.
@@
The effect of trapped charges can be illustrated by simulating channel formation in a typical QD device. Figures 1(a-b) show a gate-defined, accumulation-mode, undoped QD device with multiple current channels, six designated regions for forming QDs (under the six plunger gates, shown in red), and reservoir gates (shown in blue) to connect these regions to Ohmic contacts for current flow. This device exemplifies a commonly used fabrication strategy known as overlapping gates architecture [35, 36].
@@
The screening gates (shown in yellow) form the bottommost layer in the three-layer gate architecture, positioned closest to the gate oxide. Their primary role is to screen electric fields, thereby preventing the unintentional formation of 2DEG or 2DHG regions and electrically isolating the intended transport channels. The second layer includes the reservoir and plunger gates. The reservoirs enable the flow of electrons from the Ohmic contacts toward the center of the device, where QDs form under electrostatic confinement. Each plunger gate primarily controls the chemical potential of the QD below it. The final layer consists of barrier gates (shown in gray), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and reservoirs. By careful adjustment of the voltages applied to each gate, the QD qubit chip with charge sensors is made functional [2, 3, 4, 5].
@@
Figure 1: (a) False-color scanning electron microscopy image of a device nominally identical to one used in this work. The color indicates the gate’s function as a screening (S1-S4; yellow), reservoir (R1-R5; blue), barrier (B1-B9; gray), or plunger (P1-P6; red) gate. (b) Higher-magnification image with the desired current paths shown by black arrows and the individual barrier and plunger gates labeled. (c-f) Self-consistent Schrödinger-Poisson simulations (density plots), showing typical effects of trapped charge on 1D channel formation in a 2DEG. The simulation results are overlaid on the scanning electron microscopy image. (c) Trapped charge can lead to current paths crossing the central screening gate; an autotuner must detect and correct for this problem, as shown in (d). (e) Trapped charge can also lead to breaks in the current paths, another possible error that an autotuner must detect and correct for, as shown in (f). The insets in (c)–(f) show the gate voltages necessary to cause the charge densities. For all simulations, interface trapped charge is chosen, as described in the main text and in the Supplemental Material [37].
@@
To model the effects of trapped charge, we introduce randomly varying charge densities on a tiled grid array at the oxide-semiconductor interface, with the density of each tile selected at random from a normal distribution, with the standard deviation corresponding to a shift magnitude $\Delta V=83$ $\mathrm{mV}$, consistent with experimental observations for silicon QDs [28]. Each tile is $40$ $\mathrm{nm}$ by $50$ $\mathrm{nm}$, ensuring that all gates encompass multiple regions of randomly selected charge densities.
@@
Figure 2: BATIS algorithm. The flowchart representation of the procedural flow of BATIS. Beyond manipulating gate voltages, BATIS can also change the device response to gate voltages in situ by illuminating the device. Active steps in autotuning are shown by rounded gray rectangles, conditional logic is shown by yellow diamonds, and failure determinations are shown by red stop signs. The green and red arrows indicate whether a given step passes or fails, respectively.
@@
In the case where no random trapped charge is included, the ideal charge densities in the quantum well form coherent channels half way between the screening gates, as indicated by the arrows in Fig. 1(b). In contrast, Figs. 1(c) and 1(e) show results obtained when randomly distributed trapped charges are included, resulting in current channels that flow beneath screening gates and pinched channels, respectively. Both problems can be resolved by careful adjustment of the various gate voltages to recover the coherent channels shown in Figs. 1(d) and 1(f), respectively.
@@
While misdirected and pinched channels are easy to identify in simulations, they are much more difficult to diagnose in experiments. The issue of current flowing beneath the screening gate and unintentionally connecting multiple channels, as illustrated in Fig. 1(c), is intrinsic to 2D gate-defined systems and does not exist in one-dimensional (1D) nanowire and nanotube devices [38, 39]; therefore, it is of particular interest in this work. The recovery of misdirected current to form the desired coherent channels requires the simultaneous tuning of multiple gates. The protocols developed in this work are specially designed to address and characterize such 2D challenges.
Figure 3: Stages of the tuning procedure. The first row in (a)–(d) explains the measurements. The second row shows idealized cartoon representations of the key data for each stage of the tuning process, with good and bad data shown on the left and the right, respectively. The third row provides examples of experimental data. The yellow arrows in the cartoons indicate the ideal current paths. (a) In the leakage test, no gate should have a low resistance to any other gate or cryostat ground (Ohmic contacts are included in the test for the nonaccumulated leakage test). The good cartoon matrix shows high-resistance paths on all connections. The bad cartoon matrix shows two gates shorted to each other. The realistic matrix shows three gates shorted to ground. Devices with shorted gates are typically unusable. (b) When accumulating the 2DEG in the device globally, the current is expected to rise above zero with the increase of gate voltages, as seen in the good cartoon plot. The bad cartoon plot shows a failure to turn on. A sccessful turn-on is observed in the experimental data, with G representing $\{{\rm S}_{1},{\rm S}_{3},{\rm S}_{4},{\rm R}_{1},{\rm R}2,{\rm R}3,{\rm R}4,{\rm R}5\}$ set of gates. (c) For pinch-off, it is desirable for the current to drop to the noise floor as the gate voltage is lowered. The bad cartoon demonstrates a failure of the screening or reservoir gate to pinch off. The experimental data show some small oscillations due to, for example, accidental Coulomb blockade during pinch-off. (d) For channel accumulation, rather than relying on measuring pairs of screening gates, we sweep a single screening gate against all the finger gates corresponding to that current channel. The bad cartoon measurement indicates that at least one finger gate is too far below the threshold to accumulate. (e) Each plunger and barrier gate must pinch off individually. The bad cartoon shows the finger gate failing to pinch off. In the experimental data, Coulomb blockade oscillations are often observed and are particularly prominent during pinch-off at this tuning step.
@@ III BATIS: Design and implementation
BATIS is designed to automate the testing, characterization, and initialization of all gates, as well as the formation of all necessary 1D current channels for linear QD arrays. The procedural flow of the single-button-press algorithm, depicted in Fig. 2, is rooted in the underlying device physics, making it both interpretable and robust, while also enabling adaptability across devices and scalability to arbitrary-length 1D arrays.
@@
The configuration files establish a link between abstract descriptions of device components and the operational language used within BATIS. For instance, the functionality of each gate is automatically inferred from the user-provided device architecture, enabling BATIS to construct a gate lookup table. When combined with the wire map, this allows BATIS to automatically identify and perform the appropriate specialized measurements by linking the logical gate definitions to the physical measurement instruments. This abstraction layer ensures that BATIS remains agnostic to both the specific device layout and the instrumentation used.
@@
The algorithm follows conditional logic, which is visually represented by boxes and arrows of different colors in Fig. 2. Each of the six stages in BATIS, indicated in Fig. 2 with rounded gray rectangles, involves at least one optimization of a device measurement or manipulation followed by data analysis to extract characteristics of a selected subset of gates. The measurements typical of each stage are detailed in Fig. 3, with the first row in each panel explaining the measurement, and the second and third rows showing a cartoon representation of the data (in each case, showing success on the left and failure on the right) and an example of measured experimental data for each step, respectively.
@@
The procedural steps in BATIS are indicated by gray arrows. Passing the functionality test for the current stage, marked with yellow diamonds in Fig. 2, initiates the next stage, as indicated by the green arrows. Failing a test, indicated by red arrows, results in repeated measurements, device optical illumination to adjust the trapped charges, or bootstrapping termination, depending on the issues revealed by BATIS and the conclusions it draws on the basis of the device physics. In the present implementation, illumination is treated as a rare, high-impact recovery step rather than a continuously tunable control knob. Each illumination heats the cryostat and, therefore, only a small number of illumination attempts are feasible per cooldown. Within this very small, safety-constrained action space, a simple, physics-motivated decision tree provides a transparent and effectively sample-optimal policy for illumination, establishing a baseline against which more sophisticated controllers can later be benchmarked.
@@
When the procedures have been completed, BATIS produces a device diagnostic file that stores all information about each gate learned during the tuning process. When executed successfully, it verifies that all gates are functioning as intended and that isolated regions can be formed in well-defined 1D channels beneath the designated plunger gates. It also preconfigures the QD device for further use. In the following subsections, we explain each state of the algorithm.